# Extremal distributions of partially hyperbolic systems: the Lipschitz threshold

### 01 Apr 2026 v1

@@ 6.2. Proof of Corollary B
Let $\Lambda$ be a compact $f$-invariant set. If $\Lambda$ is $\ell$-integrable for some $\ell>2$, then $E^{s}\oplus E^{u}$ is Lipschitz on $\Lambda$.
@@
Since $\{S_{x}\}$ is a continuous family of $C^{\ell}$ surfaces with $\ell>2$, this implies that $E^{s}\oplus E^{u}$ is uniformly $C^{1}$ along $\mathcal{W}^{u}$.
By Lemma 6.4, $E^{s}\oplus E^{u}$ is uniformly $C^{1}$ along $\mathcal{W}^{c}_{\mathrm{loc}}(x)$ for every $x\in M$. Since $\mathcal{W}^{s}_{\mathrm{loc}}$, $\mathcal{W}^{u}_{\mathrm{loc}}$, and $\mathcal{W}^{c}_{\mathrm{loc}}$ give a local product structure in a neighborhood of $x$ for every $x\in M$, we deduce that $E^{s}\oplus E^{u}$ is Lipschitz [^3] Note that we cannot apply Journé’s lemma [42] to obtain $C^{1}$ regularity since $\mathcal{W}^{c}$ may not form a foliation. on $\Lambda$. ∎
@@
If $E^{s}\oplus E^{u}$ is integrable, then by Lemma 2.10 there exists an integral foliation $\mathcal{W}^{su}$ with uniformly $C^{\infty}$ leaves, and $f$ is $\ell$-integrable for any $\ell>2$.
@@ 6.3. Proof of Corollary C
First, consider case (1), i.e., when $f$ is accessible. In this case, $f$ is ergodic with respect to $m$, and since $m(\mathrm{supp}(\mu))>0$, it follows that $\mathrm{supp}(\mu)=M$. As $E^{s}\oplus E^{u}$ is not integrable, Corollary B implies that $f$ is not $\ell$-integrable for any $\ell>2$, and by [27, Theorem 8.1], $\mu$ satisfies the QNI condition. Furthermore, [27, Theorem 1.2] ensures that $\mu$ is physical, meaning its basin has positive $m$-measure (hence full measure, by ergodicity of the volume measure $m$). Applying Birkhoff’s ergodic theorem to a point in the intersection of the basins of $m$ and $\mu$, we conclude that $\mu=m$.
We now consider case (3) of Proposition 2.7. Since $m(\mathrm{supp}(\mu))>0$, Lemma 2.8 implies that $\mathrm{supp}(\mu)=\overline{U}$, where $U:=\bigcup_{k=0}^{n-1}f^{k}(U_{0})$ for some $f^{n}$-invariant open accessibility class $U_{0}$ and integer $n\geq 1$. Let us consider the $f$-invariant compact set $\Lambda:=\overline{U}=\mathrm{supp}(\mu)$. By point (2) of Corollary B, $\mathrm{supp}(\mu)=\Lambda$ is not $\ell$-integrable for any $\ell>2$; it follows from [27, Theorem 8.1] that $\mu$ satisfies the QNI condition. Furthermore, [27, Theorem 1.2] implies that $\mu$ is physical. As in case (1), Birkhoff’s ergodic theorem yields $\mu=m_{U}$, where $m_{U}:=\frac{m|_{U}}{m(U)}$. This completes the proof.
@@ 6.4. Proof of Corollary D
  • an isometric extension of an Anosov diffeomorphism of $\mathbb{T}^{2}$;
@@
In the first case, Theorem A implies that $f$ is Anosov and that the center cocycle $x\mapsto\lambda^{c}_{x}$ is cohomologous to a constant. Replacing $f$ by $f^{-1}$ if necessary, we may assume that $f$ is uniformly expanding along $\mathcal{W}^{c}$. Since $\mathcal{W}^{c}$ is absolutely continuous, it follows from [35, Proposition 2] that the unstable cocycle $x\mapsto\lambda^{u}_{x}$ is cohomologous to a constant. As $f$ preserves volume, the multiplicative cocycle $x\mapsto\lambda^{s}_{x}\lambda^{u}_{x}\lambda^{c}_{x}$ is a coboundary, and hence so is the stable cocycle $x\mapsto\lambda^{s}_{x}$. Therefore, by [24, Theorem 1.1], $f$ is $C^{\infty}$-conjugate to its linear part.
@@
In the remaining cases, Proposition 6.1 implies that $f$ is $C^{\infty}$-conjugate to the corresponding model. Moreover, if $f$ is $C^{\infty}$-conjugate to the time-one map of an Anosov flow $\varphi_{t}$ with $E^{s}\oplus E^{u}$ of class $C^{1}$, then $\varphi_{t}$ is either a contact flow or a suspension flow [52] (see also [29, Theorem 3]).

### 16 Apr 2026 v2

@@ 6.2. Proof of Corollary B
Let $\Lambda$ be a compact $f$-invariant set. If $\Lambda$ is $\ell$-integrable for some $\ell\geq 2$, then $E^{s}\oplus E^{u}$ is $C^{1}$ on $\Lambda$.
@@
Since $\{S_{x}\}$ is a continuous family of $C^{\ell}$ surfaces with $\ell\geq 2$, this implies that $E^{s}\oplus E^{u}$ is uniformly $C^{1}$ along $\mathcal{W}^{u}$.
By Lemma 6.4, the bundle $E^{s}\oplus E^{u}$ is uniformly $C^{1}$ along each center leaf $\mathcal{W}^{c}_{\mathrm{loc}}(x)$ for every $x\in M$. Since the foliations $\mathcal{W}^{s}_{\mathrm{loc}}$, $\mathcal{W}^{u}_{\mathrm{loc}}$, and $\mathcal{W}^{c}_{\mathrm{loc}}$ form a local product structure near every point and vary continuously, it follows from a standard calculus argument that $E^{s}\oplus E^{u}$ is $C^{1}$ on $\Lambda$. ∎
@@
If $E^{s}\oplus E^{u}$ is integrable, then by Lemma 2.10 there exists an integral foliation $\mathcal{W}^{su}$ with uniformly $C^{\infty}$ leaves, and $f$ is $\ell$-integrable for any $\ell$.
@@ 6.3. Proof of Corollary C
First, consider case (1), i.e., when $f$ is accessible. In this case, $f$ is ergodic with respect to $m$, and since $m(\mathrm{supp}(\mu))>0$, it follows that $\mathrm{supp}(\mu)=M$. As $E^{s}\oplus E^{u}$ is not integrable, Corollary B implies that $f$ is not $\ell$-integrable for any $\ell>2$, and by [28, Theorem 8.1], $\mu$ satisfies the QNI condition. Furthermore, [28, Theorem 1.2] ensures that $\mu$ is physical, meaning its basin has positive $m$-measure (hence full measure, by ergodicity of the volume measure $m$). Applying Birkhoff’s ergodic theorem to a point in the intersection of the basins of $m$ and $\mu$, we conclude that $\mu=m$.
We now consider case (3) of Proposition 2.7. Since $m(\mathrm{supp}(\mu))>0$, Lemma 2.8 implies that $\mathrm{supp}(\mu)=\overline{U}$, where $U:=\bigcup_{k=0}^{n-1}f^{k}(U_{0})$ for some $f^{n}$-invariant open accessibility class $U_{0}$ and integer $n\geq 1$. Let us consider the $f$-invariant compact set $\Lambda:=\overline{U}=\mathrm{supp}(\mu)$. By point (2) of Corollary B, $\mathrm{supp}(\mu)=\Lambda$ is not $\ell$-integrable for any $\ell>2$; it follows from [28, Theorem 8.1] that $\mu$ satisfies the QNI condition. Furthermore, [28, Theorem 1.2] implies that $\mu$ is physical. As in case (1), Birkhoff’s ergodic theorem yields $\mu=m_{U}$, where $m_{U}:=\frac{m|_{U}}{m(U)}$. This completes the proof.
@@ 6.4. Proof of Corollary D
  • an isometric extension of an Anosov diffeomorphism on $\mathbb{T}^{2}$;
@@
In the first case, Theorem A implies that $f$ is Anosov and that the center cocycle $x\mapsto\lambda^{c}_{x}$ is cohomologous to a constant. Replacing $f$ by $f^{-1}$ if necessary, we may assume that $f$ is uniformly expanding along $\mathcal{W}^{c}$. Since $\mathcal{W}^{c}$ is absolutely continuous, it follows from [37, Proposition 2] that the unstable cocycle $x\mapsto\lambda^{u}_{x}$ is cohomologous to a constant. As $f$ preserves volume, the multiplicative cocycle $x\mapsto\lambda^{s}_{x}\lambda^{u}_{x}\lambda^{c}_{x}$ is a coboundary, and hence so is the stable cocycle $x\mapsto\lambda^{s}_{x}$. Therefore, by [25, Theorem 1.1], $f$ is $C^{\infty}$-conjugate to its linear part.
@@
In the remaining cases, Proposition 6.1 implies that $f$ is $C^{\infty}$-conjugate to the corresponding model. Moreover, if $f$ is $C^{\infty}$-conjugate to the time-one map of an Anosov flow $\varphi_{t}$ with $E^{s}\oplus E^{u}$ of class $C^{1}$, then $\varphi_{t}$ is either a contact flow or a suspension flow [54] (see also [30, Theorem 3]).
6.5. Proof of Corollary F
Lemma 6.9.
We have $R(x)\in E^{c}(x)$ for all $x\in M$, and $\ker\alpha=E^{s}\oplus E^{u}$, where $R\subset TM$ is the Reeb vector field associated with the contact form $\alpha$.
Proof.
Since $\alpha$ is $f$-invariant, the Reeb vector field $R$ is $f$-invariant and bounded, which implies that $R(x)\in E^{c}(x)$, for all $x\in M$. For any $v\in E^{s}(x)$, we have
$\alpha(v)=((f^{n})^{\ast}\alpha)(v)=\alpha(Df^{n}(v))\to 0$
as $n\to\infty$, which implies $E^{s}\subset\ker\alpha$. Similarly, $E^{u}\subset\ker\alpha$, and hence $\ker\alpha=E^{s}\oplus E^{u}$. ∎
To apply Theorem A, it remains to show that $f$ preserves a $C^{\infty}$ volume form.
Lemma 6.10.
The diffeomorphism $f$ is accessible.
Proof.
We show that every accessibility class $A$ is open. Since $M$ is connected and accessibility classes are pairwise disjoint, this implies that $f$ is accessible.
We argue by contradiction. By [21, Proposition 5], if $A$ is not open, then $f$ is jointly integrable at every point $x\in A$, i.e., there exists $\epsilon>0$ such that for all $y\in\mathcal{W}^{u}_{\epsilon}(x)$ and $z\in\mathcal{W}^{s}_{\epsilon}(x)$, we have
$\mathcal{W}^{s}_{\mathrm{loc}}(y)\cap\mathcal{W}^{u}_{\mathrm{loc}}(z)\neq\emptyset.$
In particular, choose $w\in\mathcal{W}^{s}_{\mathrm{loc}}(y)\cap\mathcal{W}^{u}_{\mathrm{loc}}(z)$, and let $\Omega$ be a surface transverse to $E^{c}$ whose boundary consists of stable and unstable plaques joining $x,y,z,w$. Since $\ker\alpha=E^{s}\oplus E^{u}$, the integral of $\alpha$ along stable and unstable arcs (contained in leaves of $\mathcal{W}^{s}$ or $\mathcal{W}^{u}$) vanishes. Since $d\alpha$ is non-degenerate on $\Omega$, Stokes’ theorem gives
$0\neq\int_{\Omega}d\alpha=\int_{\partial\Omega}\alpha=0,$
a contradiction. ∎
Lemma 6.11.
The diffeomorphism $f$ preserves a $C^{\infty}$ volume form $\theta$.
Proof.
Let $\theta=\alpha\wedge d\alpha$, then $\theta$ is a continuous $f$-invariant 33-form. Let $m$ be the Riemannian volume form and let $\rho$ be the density function such that $\theta=\rho m$. The invariance $f^{\ast}\theta=\theta$ implies
$\log\det(Df)=\log\rho\circ f-\log\rho.$
Since $f$ preserves the Reeb vector field $R$, it is a center isometry and, in particular, it is strongly $\infty$-bunched. By Proposition 2.13, $\rho$ is $C^{\infty}$, and hence $\theta$ is a $C^{\infty}$ volume form. ∎
Since $\alpha$ is $C^{1}$, the distribution $\ker\alpha=E^{s}\oplus E^{u}$ is also $C^{1}$. Corollary F now follows from Theorem A.
6.6. Proof of Corollary E
By Theorem A, it suffices to consider the case where $f$ is either the time-one map of an Anosov flow or an isometric extension. In the flow case, the conclusion follows immediately from [36, Theorem 4.6] and [34, Theorem 1.1].
We now consider the case where $f$ is an isometric extension. Since $\mathcal{W}^{c}$ is smooth, the quotient
$\hat{f}\colon M/\mathcal{W}^{c}\to M/\mathcal{W}^{c}$
is a $C^{\infty}$ Anosov diffeomorphism. Moreover, the induced bundles $\hat{E}^{s}$ and $\hat{E}^{u}$ are $C^{1+\mathrm{Lip}}$. By [36, Corollary 3.3], the map $\hat{f}$ is $C^{\infty}$ conjugate to an affine automorphism, and $\hat{E}^{s}$ and $\hat{E}^{u}$ are $C^{\infty}$.
Since $E^{s}\oplus E^{u}$ is $C^{\infty}$ and transverse to $E^{c}$, the projection $\pi\colon M\to M/\mathcal{W}^{c}$ induces a $C^{\infty}$ vector bundle isomorphism
$d\pi|_{E^{s}\oplus E^{u}}\colon E^{s}\oplus E^{u}\to T(M/\mathcal{W}^{c}).$
Let $\tilde{E}^{s}:=(d\pi|_{E^{s}\oplus E^{u}})^{-1}(\hat{E}^{s})$, then $\tilde{E}^{s}$ is a $C^{\infty}$ subbundle of $TM$, and it is invariant under $Df$ and exponentially contracted. By uniqueness of the stable bundle, it follows that $E^{s}=\tilde{E}^{s}$, hence $E^{s}$ is $C^{\infty}$. The same argument applies to $E^{u}$.
Finally, since $f$ is isometric along the center direction and descends to an affine automorphism on ${\mathbb{T}}^{2}$,we can choose a $C^{\infty}$ global frame $X_{\ast}\subset E^{\ast}$, $\ast=s,c,u$, such that $Df(x)$ is represented by a constant matrix for all $x\in{\mathbb{T}}^{3}$. The conclusion then follows from [1, Theorem 1.1].
6.7. An example on Heisenberg manifold
We present an example of a contact partially hyperbolic diffeomorphism on a Heisenberg manifold which is not $C^{1}$-conjugate to any affine automorphism.
Example 6.12.
Let $M=\mathbb{R}^{3}/\sim$ be the $3$-dimensional Heisenberg manifold, where
$(x,y,z)\sim(x+m,\,y+n,\,z+k+my+\tfrac{1}{2}mn),\quad m,n,k\in\mathbb{Z}.$
Let $\alpha=dz-x\,dy$ be the standard contact form on $M$. Define
$L(x,y,z)=\left(2x+y,\,x+y,\,z+x^{2}+xy+\tfrac{1}{2}y^{2}\right).$
A direct computation shows that $L^{\ast}\alpha=\alpha$.
For $\varepsilon\in\mathbb{R}$ sufficiently small, define
$u(x)=\varepsilon\sin(2\pi x),\qquad U(x)=\varepsilon x\sin(2\pi x)+\frac{\varepsilon}{2\pi}\bigl(\cos(2\pi x)-1\bigr),$
and set
$H_{\varepsilon}(x,y,z)=\bigl(x,\,y+u(x),\,z+U(x)\bigr).$
Since $U^{\prime}(x)=xu^{\prime}(x)$, we have
$H_{\varepsilon}^{\ast}\alpha=d(z+U)-x\,d(y+u)=dz-x\,dy=\alpha.$
Thus $H_{\varepsilon}$ preserves the contact form.
We now verify that $H_{\varepsilon}$ is well-defined on $M$. Using the identities
$u(x+1)=u(x),\qquad U(x+1)-U(x)=u(x),$
by a direct calculation, we can check that $H_{\varepsilon}$ commutes with the deck transformations defining the quotient. Therefore, $H_{\varepsilon}$ descends to a well-defined diffeomorphism on $M$.
Define
$F_{\varepsilon}:=L\circ H_{\varepsilon}.$
Then $F_{\varepsilon}$ is a contact partially hyperbolic diffeomorphism preserving the volume form $\alpha\wedge d\alpha$. By Corollary F, the distributions $E^{s}\oplus E^{u}$ and $E^{c}$ are $C^{\infty}$.
Finally, note that $\operatorname{tr}(DL(x,y,z))=4$ for all $(x,y,z)\in M$, while $F_{\varepsilon}(0)=0$ and
$\operatorname{tr}(DF_{\varepsilon}(0))=4+2\pi\varepsilon.$
It follows that $F_{\varepsilon}$ is not $C^{1}$-conjugate to $L$.
Remark 6.13.
In the example above, the map $F_{\epsilon}$ is volume-preserving, and the distributions $E^{s}\oplus E^{u}$ and $E^{c}$ are $C^{\infty}$. For any $\delta>0$, by choosing $\epsilon=\epsilon(\delta)$ sufficiently small, the distributions $E^{s}$ and $E^{u}$ are $C^{2-\delta}$. However, the distributions $E^{s}$ and $E^{u}$ are not $C^{1+\mathrm{Lip}}$ by Corollary E.